Investments

Chaplain Company was very active in acquiring and selling investments in equity securities. Data regarding the securities are:


December 31, 2020                                 COST          MARKETVALUE

Trading Securities                                   5,000,000         4,600,000

Securities not held for trading                   3,000,000         3,100,000


December 31, 2021

Trading Securities                                    5,000,000         5,500,000

Securities not held for trading                  3,000,000            3,300,000

   

The entity made an irrevocable election to present changes in fair value of the securities not held for trading in other comprehensive income.

1. How much is the unrealized gain/loss in trading recognized in P2020? 

2. How much is the unrealized gain/loss to be recognized P2021?


Prepare the journal entries to recognized the changes in market value for 2020 and 2021.

3. Entry in 2020 ?

 

4. Entry in 2021 ?

 

5. Assuming the Trading Securities was sold at 6,000,000 on December 31, 2021, How much was the realized gain/loss on sale

a. 500,000 loss

b. 500,000 gain

c. 1,000,000 gain

d. 1,000,000 loss
Answer #1

1) Unrealized gain/loss in trading recognized in P2020

The loss to be recognized in Profit and Loss Account= Market Value of Trading Securities- Cost of Trading Securities

=4600000-5000000=(400000)

2)Unrealized gain/loss to be recognized P2021

Gain to be recognized in Profit and Loss Account= Market Value of Trading Securities- Cost of Trading Securities

=5500000-4600000

=$900000

Gain to be recognized in Other Comprehensive Income=Market Value of Non-Trading Securities-Cost of Non-Trading Securities

=3300000-3000000

=$300000

Journal entries to recognize the changes in market value for 2020 and 2021.

DateParticularsDebitCredit
Dec-31 2020Profit and Loss A/c40000

To Trading Securities
40000

(Being fair value recorded)


Non-Trading Securities100000

To Other Comprehensive A/c
100000

(Being fair value recorded)

Dec-31 2021Trading Securities90000

To Profit and Loss A/c
90000

(Being fair value recorded)


Non-Trading Securities200000

To Other Comprehensive A/c
200000

(Being fair value recorded)

Realised Gain on Sale of Trading Securities= Sale Value- Market Value=6000000-5500000 =500000

55,00,000 is the Market Value as recorded in books


answered by: tutor G.
Answer #2

Entry in 2020

Securities not held for trading (FVTOCI) dr. 3100000

To Securities not held for trading (AC) 3000000

To OCI ( gain) 100000

Entry in 2021

Securities not held for trading (FVTOCI) dr. 200000

   To OCI ( gain) 200000


answered by: tutor G.